What is the mass of 4.5 moles of magnesium sulfate?       
Answer #1

No of moles = mass/molecularmass

mass = No of moles * molarmass

Molarmass of MgSO4 = 120.3676 g/mol

No of moles = 4.5

Mass of MgSO4= 4.5*120.3676 = 541.6542 grams
Answer #2

No of moles = mass/molecularmass

mass = No of moles * molarmass

Molarmass of MgSO4= 120.366 g/mol

No of moles = 4.5

Mass of NaI = 4.5*120.366 =541.647 grams
Answer #3

No of moles = mass/molecularmass

mass = No of moles * molarmass

Molarmass of MgSO4 = 120.366 g/mol

No of moles = 4.5

Mass of NaI = 4.5*120.366 = 541.647 grams